I am so sorry. This is the problem
Solve each equation. Identify each equation as a conditional equation, an inconsistent equation, or an identity
1/2b - 1/2 = 1/4b
Please help me with this problem. I wrote it down wrong the last time. thank so much

Answer by fractalier(6550)

I did this one already, but here goes again...from
1/2b - 1/2 = 1/4b
we multiply all by the lowest common denominator, 4b, and get
4b[1/2b - 1/2 = 1/4b]
2 - 2b = 1
-2b = -1
b = 1/2